OpenStudy (ashking1):

Use the arc length formula and the given information to find s. r = 20 ft θ = 19° s = ?

OpenStudy (mathstudent55):

\(\Large s = \dfrac{\theta}{360^o}2 \pi r\) where s = arc length \(\theta\), Greek letter "theta" (pronounced "thayta" or "theeta"), = angle measure r = radius of circle